A and B decided to meet between 3 PM, and 4 PM, each of them waits only 10 minutes. what is the probability that they will meet?


1
Expert's answer
2021-01-19T04:07:18-0500
"|x-y|\leq\dfrac{10}{60}"

"P=1\times1-2(\dfrac{1}{2}\times(1-\dfrac{1}{6})\times(1-\dfrac{1}{6}))=\dfrac{11}{36}"

The probability that they will meet is "\dfrac{11}{36} ."